an investor puts £5000 in a savings account that pays 10% simple interest at the end of each year. compare how much the investor would have after 6 years if the money was invested for 6 years?


1
Expert's answer
2021-04-13T16:08:29-0400

If invested

"A= P(1+rt)"

where A = new balance

t= time

r= rate in decimals

P = Amount invested

"A= 5000 (1+01*6)"

"=5000(1.6)"

"= 8000"

£8000


In a Savings account

"A=P(1+r\/n)^n*^t"

n= number of times its compounded

"A=5000(1+0.1\/1)^1*^6"

"A=5000(1.1)^6"

"= 8857.805"

£8858

The investor will have £858 more if he saves